A particle of mass 1kg is moving along the line y=x+2 with speed 2m/s. The magnitude of angular momentum of particle about origin is??

Grade:11

2 Answers

Aman Bansal
12 years ago

Dear Sai,

y = x + 2

r = x i, x + 2 j
v = sqrt(2) * (i + j)

angular momentum = r X mv = 1 * sqrt(2){(x i + (x + 2) j ) X (i + j) } = sqrt(2) (x k - x + 2k) = -2 sqrt(2) k

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert


Sakshi Bhavik
5 years ago
Given data : Mass of the particles = 1 kg
Velocity of the particle = 2 m/s
Given Equation of motion of the particle : y = x+4 which indicates a straight line.
So , Angular momentum of the particle about the origin, L = r X p = mvr  sin θ where r is the perpendicular distance between the line and the origin.  Here, sin  θ = 1.
Thus , Angular momentum of the particle about the origin = mvr = 1 kg * 2m/s * 2m = 4 kg.m2/s